# spacing=nonfrench

### v1

@@ 5.4 Improvements through Weight Averaging
[…]
@@ 6.1 Zero-Shot Adaptive Compute at Test-Time
We show this zero-shot per-token adaptive compute behavior in Figure 10, where we plot the distribution of steps taken before the exit condition is hit. We do this for the first 50 questions from different MMLU categories, asked in free-form chat. Interestingly, the number of steps required to exit differs notably between categories, with the model exiting earlier on high school mathematics, but taking on average 3.5 steps more on moral scenarios. As a preliminary demonstration, we verify on MTBench that this adaptivity does not significantly impact performance in a conversational benchmark setting (standard: 5.635.635.63, early exits: 5.565.565.56 see Appendix Table 6).
@@
Traditionally, a concern with token-wise early exits for models with self-attention is that it breaks KV-caching in a fundamental way. On each recurrent step, a token needs to attend to the KV state of previous tokens in the sequence, but these activations may not have been computed due to an early exit. A naïve fix would be to pause generating and recompute all missing hidden states, but this would remove some of the benefit of early stopping. Instead, as in Elbayad et al. (2019), we attend to the last, deepest available KV states in the cache. Because all recurrent KV cache entries are generated by the same K,V projection matrices from successive hidden states, they ``match'', and therefore the model is able to attend to the latest cache entry from every previous token, even if computed at different recurrent depths.
@@ 6.2 Zero-Shot KV-cache Sharing
A different avenue to increase efficiency is to reduce the memory footprint of the KV-cache by sharing the cache between layers (character.ai, 2024; Brandon et al., 2024). Typically, transformers must be trained from scratch with this capability. However, as discussed in the previous section, we find that we can simply share KV-caches in our model with minimal impact to performance. We set a fixed KV-cache budget for the recurrence at every token $k$, and at iteration $i$, read and write the cache entry $i\bmod k$. For example, we set a maximum KV-cache budget of 16 steps, overwriting the KV-cache of the 1st step when executing the 17th step, and so forth. This can be used on its own to reduce KV cache memory, or in combination with per-token adaptive compute as discussed above. On MTBench, this does not reduce performance (cache budget of 444: 5.865.865.86, see Appendix Table 6).
Figure 11: Convergence of latent states for every token in a sequence (going top to bottom) and latent iterations (going left to right), plotting the distance a final iterate $s^{*}$, which we set with $r=128$. Shown is an unsafe question posed to the model. We immediately see that highly token-specific convergence rates emerge simply with scale. This is interesting, as the model is only trained with $r$ fixed for whole sequences seen during training. We see that convergence is especially slow on the key part of the question, really wrong-ed.We further see that the model also learns different behaviors, we see an oscillating pattern in latent space, here most notably for the school token.
@@ 6.3 Zero-Shot Continuous Chain-of-Thought
Instead of sampling a random initial state $\mathbf{s}_{0}$ at every generation step, we can warm-start with the last state $\mathbf{s}_{r}$ from the previous token. As shown in Figure 10, this reduces the average number of steps required to converge by 1-2. Also, on tasks such as philosophy questions, we see that the exit distribution shifts on several tasks, with the model more often exiting early by recycling previous compute. To achieve a similar behavior in fixed-depth transformers, these models need to be trained on reasoning tasks to accept their last hidden state as alternative inputs when computing the next token (Hao et al., 2024).
@@ 6.4 Zero-Shot Self-Speculative Decoding
Figure 12: Latent Space trajectories for select tokens. We show a small part of these high-dimensional trajectories by visualizing the first 6 PCA directions, computing the PCA over all latent state trajectories of all tokens in a sequence. The color gradient going from dark to bright represents steps in the trajectory. The center of mass is marked in red. While on many tokens, the state simply converges (top row), the model also learns to use orbits (middle row), and ``sliders'' (bottom row, middle) to represent and handle more advanced concepts, such as arithmetic or complicated deliberation.
@@ 7 What Mechanisms Emerge at Scale in Recurrent-Depth Models
Figure 11 shows the norm distance $||\mathbf{s}_{i}-\mathbf{s}^{*}||$ between each $\mathbf{s}_{i}$ in a trajectory and an approximate limit point $\mathbf{s}^{*}$ computed with 128 iterations. We show the sentence top to bottom and iterations from left to right. We clearly see that convergence behavior depends on context. We see that key parts of the question, and the start of the model response, are ``deliberated'' much more in latent space. The context dependence can also be seen in the different behavior among the three identical tokens representing each of the three dots. Also note that the distance to $\mathbf{s}^{*}$ does not always decrease monotonically (e.g. for school); the model may also trace out complicated orbits in its latent trajectory while processing information, even though this is not represented explicitly in our training objective.
We look at trajectories for select tokens in more detail in Figure 12. We compute a PCA decomposition of latent trajectories over all tokens in a sequence, and then show several individual trajectories projected onto the first six PCA directions. See the appendix for more examples. Many tokens simply converge to a fixed point, such as the token in the top row. Yet, for harder questions, such as in the 2nd row [^4] This is the token ”3” in a GSM8k test question that opens with Claire makes a 3 egg omelette., the state of the token quickly falls into an orbit pattern in all three pairs of PCA directions. The use of multi-dimensional orbits like these could serve a similar purpose to periodic patterns sometimes observed in fixed-depth transformers trained for arithmetic tasks (Nanda et al., 2022), but we find these patterns extend far beyond arithmetic for our model. We often observe use of orbits on tokens such as ``makes'' (see Figure 16) or ``thinks'' that determine the structure of the response.
Aside from orbits, we also observe the model encoding particular key tokens as ``sliders'', as seen in the middle of the bottom row in Figure 12 (which is the token ``wrong'', from the same message as already shown in Figure 11). In these motions the trajectory noticeably drifts in a single direction, which the model could use to implement a mechanism to count how many iterations have occurred.
@@
We verify that our models maintain path independence, in the sense of Anil et al. (2022), despite their complex dynamics, which we demonstrate in Figure 22. When re-initializing from multiple starting points $\mathbf{s}_{0}$, the model moves in similar trajectories, exhibiting consistent behavior. The same orbital patterns, fixed points, or directional drifts emerge regardless of initialization.
@@ 8 Related Work Overview
The extent to which recurrence is a foundational concept of machine learning is hard to overstate (Braitenberg, 1986; Gers & Schmidhuber, 2000; Sutskever et al., 2008). For transformers, recurrence was applied in Dehghani et al. (2019), who highlight the aim of recurrent depth to model universal, i.e. Turing-complete, machines. It was used at scale (but with fixed recurrence) in Lan et al. (2019) and an interesting recent improvement in this line of work are described in Tan et al. (2023); Abnar et al. (2023) and Csordás et al. (2024). Schwarzschild et al. (2021b); Bansal et al. (2022); Bear et al. (2024); McLeish et al. (2024) show that depth recurrence is advantageous when learning generalizable algorithms when training with randomized unrolling and input injections. Recent work has described depth-recurrent, looped, transformers and studied their potential benefits with careful theoretical and small-scale analysis (Giannou et al., 2023; Gatmiry et al., 2024; Yang et al., 2024a; Fan et al., 2025).
@@
From another angle, these models can be described as neural networks learning a fixed-point iteration, as studied in deep equilibrium models (Bai et al., 2019; 2022). They are further related to diffusion models (Song & Ermon, 2019), especially latent diffusion models (Rombach et al., 2022), but we note that language diffusion models are usually run with a per-sequence, instead of a per-token, iteration count (Lee et al., 2018). A key difference of our approach to both equilibrium models and diffusion models is in the training objective, where equilibrium methods solve the ``direct'' problem (Geiping & Moeller, 2019), diffusion models solve a surrogate training objective, and our work suggests that truncated unrolling is a scalable alternative.
@@
Mor generally, architectures that recur in depth can also be understood as directly learning the analog to the gradient of a latent energy-based model (LeCun & Huang, 2005; LeCun, 2022), or to an implicitly defined intermediate layer (Amos & Kolter, 2017). These analogies to gradient descent at inference time also show the connection to test time adaptation (Sun et al., 2020), especially test-time adaptation of output states (Boudiaf et al., 2022).
@@
Aside from full recurrent-depth architectures, there also exist a number of proposals for hybrid architectures, such as models with latent subnetworks (Li et al., 2020a), LoRA adapters on top of weight-shared layers (Bae et al., 2024), or (dynamic) weight-tying of trained models (Hay & Wolf, 2023; Liu et al., 2024). As mentioned in Section 6, while we consider the proposed recurrent depth approach to be a very natural way to reason in continuous latent space, the work of Hao et al. (2024) discusses how to finetune existing fixed-depth transformers with this capability.
@@ 9 Future Work
Another aspect not covered in this work is the relationship to other modern architecture improvements. Efficient sequence mixing operations, especially those that are linear in sequence dimension, such as linear attention (Katharopoulos et al., 2020; Yang et al., 2024b), are limited in the number of comparisons that can be made. However, with recurrent depth, blocks containing linear operators can repeat until all necessary comparisons between sequence elements are computed (Suzgun et al., 2019). For simplicity, we also focus on a single recurrence, where prior work has considered multiple successive recurrent stages (Takase & Kiyono, 2023; Csordás et al., 2024).
Finally, the proposed architecture is set up to be compute-heavy, with more ``materialized'' parameters than there are actual parameters. This naturally mirrors mixture-of-expert models (MoE), which are parameter-heavy, using fewer active parameters per forward pass than exist within the model (Shazeer et al., 2017; Fedus et al., 2022). We posit that where the recurrent-depth setup excels at learning reasoning patterns, the MoE excels at effectively storing and retrieving complex information. Their complementarity supports the hypothesis that a future architecture would contain both modifications. While in a standard MoE model, each expert can only be activated once per forward pass, or skipped entirely, a recurrent MoE model could also refine its latent state over multiple iterations, potentially routing to the same expert multiple times, before switching to a different one (Tan et al., 2023; Csordás et al., 2024). While MoE models are the currently leading solution to implement this type of ``memory'' in dense transformers, these considerations also hold for other memory mechanisms suggested for LLMs (Sukhbaatar et al., 2019; Fan et al., 2021; Wu et al., 2022; He et al., 2024).
@@ Acknowledgements
This project was made possible by the INCITE program: An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. JG further acknowledges the support of the Hector II foundation. A large number of small-scale and preliminary experiments were made possible through the support of the Intelligent Systems compute cluster and funding by the Tübingen AI center. UMD researchers were further supported by the ONR MURI program, DARPA TIAMAT, the National Science Foundation (IIS-2212182), and the NSF TRAILS Institute (2229885). Commercial support was provided by Capital One Bank, the Amazon Research Award program, and Open Philanthropy. Finally, we thank Avi Schwarzschild for helpful comments on the initial draft.
@@ Appendix A Additional Information
Figure 14: Multi-Operand Arithmetic. Following a precedent of training recurrent architectures for algorithmic and arithmetic tasks (Schwarzschild et al., 2021b; Bansal et al., 2022; Schwarzschild et al., 2023; McLeish et al., 2024), we explore whether our model can leverage increased test-time compute via recurrence to solve verbalized addition problems of increased difficulty. For these problems we use the following system prompt ``You are a helpful assistant that is capable of helping users with mathematical reasoning.'' embedded in a conversational chat template, and we present each problem by opening the first user turn of the conversation like so: f"What is the result of ' + '.join(map(str, digits))?" after randomly sampling numbers according to a certain operand count and digit count (base 10). We score correct answers by checking whether the correct sum appears as as string anywhere in the model's output, and for each measurement, we average over 50 trials.

### v2

@@ 5.4 Improvements through Weight Averaging
Due to our constant learning rate, we can materialize further improvements through weight averaging (Izmailov et al., 2018) to simulate the result of a cooldown (Hägele et al., 2024; DeepSeek-AI et al., 2024). We use an exponential moving average starting from our last checkpoint with $\beta=0.9$, incorporating the last 75 checkpoints with a dilation factor of 7, a modification to established protocols (Kaddour, 2022; Sanyal et al., 2024). We provide this EMA model as well, which further improves GMS8k performance to $47.23\%$ flexible ($38.59\%$ strict), when tested at $r=64$.
@@ 6.1 Zero-Shot Adaptive Compute at Test-Time
We show this zero-shot per-token adaptive compute behavior in Figure 10, where we plot the distribution of steps taken before the exit condition is hit. We do this for the first 50 questions from different MMLU categories, asked in free-form chat. Interestingly, the number of steps required to exit differs notably between categories, with the model exiting earlier on high school mathematics, but taking on average 3.5 steps more on moral scenarios. As a preliminary demonstration, we verify on MTBench that this adaptivity does not significantly impact performance in a conversational benchmark setting (standard: $5.63$, early exits: $5.56$ see Appendix Table 6).
@@
Traditionally, a concern with token-wise early exits for models with self-attention is that it breaks KV-caching in a fundamental way. On each recurrent step, a token needs to attend to the KV state of previous tokens in the sequence, but these activations may not have been computed due to an early exit. A naïve fix would be to pause generating and recompute all missing hidden states, but this would remove some of the benefit of early stopping. Instead, as in Elbayad et al. (2019), we attend to the last, deepest available KV states in the cache. Because all recurrent KV cache entries are generated by the same K,V projection matrices from successive hidden states, they “match”, and therefore the model is able to attend to the latest cache entry from every previous token, even if computed at different recurrent depths.
@@ 6.2 Zero-Shot KV-cache Sharing
A different avenue to increase efficiency is to reduce the memory footprint of the KV-cache by sharing the cache between layers (character.ai, 2024; Brandon et al., 2024). Typically, transformers must be trained from scratch with this capability. However, as discussed in the previous section, we find that we can simply share KV-caches in our model with minimal impact to performance. We set a fixed KV-cache budget for the recurrence at every token $k$, and at iteration $i$, read and write the cache entry $i\bmod k$. For example, we set a maximum KV-cache budget of 16 steps, overwriting the KV-cache of the 1st step when executing the 17th step, and so forth. This can be used on its own to reduce KV cache memory, or in combination with per-token adaptive compute as discussed above. On MTBench, this does not reduce performance (cache budget of $4$: $5.86$, see Appendix Table 6).
Figure 11: Convergence of latent states for every token in a sequence (going top to bottom) and latent iterations (going left to right), plotting the distance a final iterate $s^{*}$, which we set with $r=128$. Shown is an unsafe question posed to the model. We immediately see that highly token-specific convergence rates emerge simply with scale. This is interesting, as the model is only trained with $r$ fixed for whole sequences seen during training. We see that convergence is especially slow on the key part of the question, really wrong-ed.We further see that the model also learns different behaviors, we see an oscillating pattern in latent space, here most notably for the school token. Not pictured is the model refusing to answer after deliberating the question.
@@ 6.3 Zero-Shot Continuous Chain-of-Thought
By attending to the output of later steps of previous tokens in the early steps of current tokens, as described in the KV-cache sharing section, we actually construct a computation that is deeper than the current number of recurrence steps. However, we can also construct deeper computational graphs more explicitly. Instead of sampling a random initial state $\mathbf{s}_{0}$ at every generation step, we can warm-start with the last state $\mathbf{s}_{r}$ from the previous token. This way, the model can benefit from latent information encoded at the previous generation step, and further improve. As shown in Figure 10, this reduces the average number of steps required to converge by 1-2. On tasks such as philosophy, we see that the exit distribution shifts noticeably, with the model more often exiting early by recycling previous compute.
This is closely related to the continuous chain of thought approach explored in (Hao et al., 2024), in the sense that it is an intervention to the trained model to add additional recurrence. To achieve a similar behavior in fixed-depth transformers, Hao et al. (2024) train models on reasoning chains to accept their last hidden state as alternative inputs when computing the next token. Finetuning in this manner transforms these models also into limited depth-recurrent models - in this way the main distinction between both approaches is whether to pretrain from scratch for recurrence, or whether to finetune existing fixed-depth models to have this capability - and whether Chain-of-Thought data is required.
@@ 6.4 Zero-Shot Self-Speculative Decoding
Figure 12: Latent Space trajectories for select tokens. We show a small part of these high-dimensional trajectories by visualizing the first 6 PCA directions, computing the PCA over all latent state trajectories of all tokens in a sequence. The color gradient going from dark to bright represents steps in the trajectory. The center of mass is marked in red. While on many tokens, the state simply converges (top row), the model also learns to use orbits (middle row), and “sliders” (bottom row, middle), which we observe being used to represent and handle more advanced concepts, such as arithmetic or complicated deliberation.
@@ 7 What Mechanisms Emerge at Scale in Recurrent-Depth Models
Figure 11 shows the norm distance $||\mathbf{s}_{i}-\mathbf{s}^{*}||$ between each $\mathbf{s}_{i}$ in a trajectory and an approximate limit point $\mathbf{s}^{*}$ computed with 128 iterations. We show the sentence top to bottom and iterations from left to right. We clearly see that convergence behavior depends on context. We see that key parts of the question, and the start of the model response, are “deliberated” much more in latent space. The context dependence can also be seen in the different behavior among the three identical tokens representing each of the three dots. Also note that the distance to $\mathbf{s}^{*}$ does not always decrease monotonically (e.g. for school); the model may also trace out complicated orbits in its latent trajectory while processing information, even though this is not represented explicitly in our training objective.
We look at trajectories for select tokens in more detail in Figure 12. We compute a PCA decomposition of latent trajectories over all tokens in a sequence, and then show several individual trajectories projected onto the first six PCA directions. See the appendix for more examples. Many tokens simply converge to a fixed point, such as the token in the top row. Yet, for harder questions, such as in the 2nd row [^5] This is the token ”3” in a GSM8k test question that opens with Claire makes a 3 egg omelette., the state of the token quickly falls into an orbit pattern in all three pairs of PCA directions. The use of multi-dimensional orbits like these could serve a similar purpose to periodic patterns sometimes observed in fixed-depth transformers trained for arithmetic tasks (Nanda et al., 2022), but we find these patterns extend far beyond arithmetic for our model. We often also observe the use of orbits on tokens such as “makes” (see Figure 16) or “thinks” that determine the structure of the response.
Aside from orbits, we also observe the model encoding particular key tokens as “sliders”, as seen in the middle of the bottom row in Figure 12 (which is the token “wrong”, from the same message as already shown in Figure 11). In these motions the trajectory noticeably drifts in a single direction, which the model could use to implement a mechanism to count how many iterations have occurred.
@@
We verify that our models maintain path independence, in the sense of Anil et al. (2022), despite their complex, learned dynamics, which we discussed prior (see also the additional examples in Appendix Figure 22). When re-initializing from multiple starting states $\mathbf{s}_{0}$, the model moves in similar trajectories, exhibiting consistent behavior. The same orbital patterns, fixed points, or directional drifts emerge regardless of initialization.
@@ 8 Related Work Overview
The extent to which recurrence is a foundational concept of machine learning is hard to overstate (Amari, 1972; Hopfield, 1982; Braitenberg, 1986; Gers and Schmidhuber, 2000; Sutskever et al., 2008). Aside from using recurrence to move along sequences, as in recurrent neural networks, it was understood early to also be the key to adaptive computation (Schmidhuber, 2012; Graves, 2017). For transformers, recurrence was applied in Dehghani et al. (2019), who highlight the aim of recurrent depth to model universal, i.e. Turing-complete, machines (Graves et al., 2014). It was used at scale (but with fixed recurrence) in Lan et al. (2019) and an interesting recent improvement in this line of work are described in Tan et al. (2023); Abnar et al. (2023), Mathur et al. (2024) and Csordás et al. (2024). Schwarzschild et al. (2021b); Bansal et al. (2022); Bear et al. (2024) and McLeish et al. (2024) show that depth recurrence is advantageous when learning generalizable algorithms when training with randomized unrolling and input injections. Recent work has described depth-recurrent, looped, transformers and studied their potential benefits with careful theoretical and small-scale analysis (Giannou et al., 2023; Gatmiry et al., 2024; Yang et al., 2024a; Fan et al., 2025).
@@
From another angle, these models can be described as neural networks learning a fixed-point iteration, as studied in deep equilibrium models (Bai et al., 2019; 2022). They are further related to diffusion models (Song and Ermon, 2019), especially latent diffusion models (Rombach et al., 2022), but we note that language diffusion models are usually run with a per-sequence, instead of a per-token, iteration count (Lee et al., 2018). A key difference of our approach to both equilibrium models and diffusion models is in the training objective, where equilibrium methods solve the “direct” problem (Geiping and Moeller, 2019), diffusion models solve a surrogate training objective, and our work suggests that truncated unrolling is a scalable alternative.
@@
More generally, all architectures that recur in depth can also be understood as directly learning the analog to the gradient of a latent energy-based model (LeCun and Huang, 2005; LeCun, 2022), to an implicitly defined intermediate optimization layer (Amos and Kolter, 2017), or to a Kuramoto layer (Miyato et al., 2024). Analogies to gradient descent at inference time also show the connection to test time adaptation (Sun et al., 2020), especially test-time adaptation of output states (Boudiaf et al., 2022).
@@
Aside from full recurrent-depth architectures, there also exist a number of proposals for hybrid architectures, such as models with latent sub-networks (Li et al., 2020a), LoRA adapters on top of weight-shared layers (Bae et al., 2024), or (dynamic) weight-tying of trained models (Hay and Wolf, 2023; Liu et al., 2024b).
As mentioned in Section 6, while we consider the proposed recurrent depth approach to be a very natural way to learn to reason in continuous latent space from the ground up, the works of Hao et al. (2024); Cheng and Durme (2024) and Liu et al. (2024a) discuss how to finetune existing fixed-depth transformers with this capability. These works have a similar aim to ours, enabling reasoning in latent space, but approach this goal from separate directions.
@@ 9 Future Work
Another aspect not covered in this work is the relationship to other modern architecture improvements. Efficient sequence mixing operations, especially those that are linear in sequence dimension, such as linear attention (Katharopoulos et al., 2020; Yang et al., 2024b), are limited in the number of comparisons that can be made. However, with recurrent depth, blocks containing linear operators can repeat until all necessary comparisons between sequence elements are computed (Suzgun et al., 2019). For simplicity, we also focus on a single recurrence, where prior work has considered multiple successive recurrent stages (Takase and Kiyono, 2023; Csordás et al., 2024).
Finally, the proposed architecture is set up to be compute-heavy, with more “materialized” parameters than there are actual parameters. This naturally mirrors mixture-of-expert models (MoE), which are parameter-heavy, using fewer active parameters per forward pass than exist within the model (Shazeer et al., 2017; Fedus et al., 2022). We posit that where the recurrent-depth setup excels at learning reasoning patterns, the MoE excels at effectively storing and retrieving complex information. Their complementarity supports the hypothesis that a future architecture would contain both modifications. While in a standard MoE model, each expert can only be activated once per forward pass, or skipped entirely, a recurrent MoE model could also refine its latent state over multiple iterations, potentially routing to the same expert multiple times, before switching to a different one (Tan et al., 2023; Csordás et al., 2024). While MoE models are the currently leading solution to implement this type of “memory” in dense transformers, these considerations also hold for other memory mechanisms suggested for LLMs (Sukhbaatar et al., 2019; Fan et al., 2021; Wu et al., 2022; He et al., 2024).
@@ Acknowledgements
This project was made possible by the INCITE program: An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Work on the LLNL side was prepared by LLNL under Contract DE-AC52-07NA27344 and supported by the LLNL-LDRD Program under Project No. 24-ERD-010 and 24-ERD-058 (LLNL-CONF-872390). This manuscript has been authored by Lawrence Livermore National Security, LLC under Contract No. DE-AC52-07NA27344 with the U.S. Department of Energy. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
JG further acknowledges the support of the Hector II foundation. A large number of small-scale and preliminary experiments were made possible through the support of the MPI Intelligent Systems compute cluster and funding by the Tübingen AI center.
UMD researchers were further supported by the ONR MURI program, DARPA TIAMAT, the National Science Foundation (IIS-2212182), and the NSF TRAILS Institute (2229885). Commercial support was provided by Capital One Bank, the Amazon Research Award program, and Open Philanthropy. Finally, we thank Avi Schwarzschild for helpful comments on the initial draft.
@@ Appendix A Additional Information
Figure 14: Multi-Operand Arithmetic. Following a precedent of training recurrent architectures for algorithmic and arithmetic tasks (Schwarzschild et al., 2021b; Bansal et al., 2022; Schwarzschild et al., 2023; McLeish et al., 2024), we explore whether our model can leverage increased test-time compute via recurrence to solve verbalized addition problems of increased difficulty. For these problems we use the following system prompt ‘‘You are a helpful assistant that is capable of helping users with mathematical reasoning.’’ embedded in a conversational chat template, and we present each problem by opening the first user turn of the conversation like so: f"What is the result of ’ + ’.join(map(str, digits))?" after randomly sampling numbers according to a certain operand count and digit count (base 10). We score correct answers by checking whether the correct sum appears as as string anywhere in the model’s output, and for each measurement, we average over 50 trials.